An average egg production of a Flock is 200 eggs per year, Sire birds with average breeding value of 300 eggs were Selected. (Assume as if Heritability of egg production and generation interval is 0.3 and eight months respectively. For a given Flock of Birds determine : A) Genetic gain per generation.

B) Genetic Gain per year.

C) Average egg Production in next Generation.


1
Expert's answer
2021-07-07T17:40:01-0400

300×3

= 900

8/12× 300 = 200

= 900+200 = 1100

1100-300 = 800 gains